OpenStudy (anonymous):

how do you solve this composition functions problem? h(x)= 3x^2-2, g(x)=3x-3.

OpenStudy (chriss):

Which composition are you looking for? g o h would be g(h(x)) and you would plug the h(x) expression in for x on the g(x) expression. \[g(h(x))=3(3x ^{2}-2)-3\] \[=9x ^{2}-6-3\] \[=9x ^{2}-9\]

OpenStudy (anonymous):

is this answer simplified?

OpenStudy (chriss):

well you could factor out the 9 to get \[9(x ^{2}-1)\]

OpenStudy (anonymous):

how would you do it if you had to foil it?

OpenStudy (chriss):

also, you didn't specify which composition, so I just picked one... you could also do a composition for h o g

OpenStudy (anonymous):

h o g is the composition. sorry.

OpenStudy (chriss):

I'm not sure what you would want to foil... although the h o g composition would require use of foil... ah.. ok.. gimme a sec to work it out.

OpenStudy (anonymous):

ok thank you.

OpenStudy (chriss):

So you start out with h(3x-3). Which means you are plugging the expression 3x-3 in for x in the expression 3x^2-2 \[h(g(x))=3(3x-3)^{2}-2\] (foil the (3x-3)^2 first) (3x-3)(3x-3) \[=9x ^{2}-18x+9\] (plug that back into the h(g(x)) expression \[3(9x ^{2}-18x+9)-2\] \[27x ^{2}-54x+27-2\] \[27x ^{2}-54x+25\]

OpenStudy (chriss):

And I think that's as simplified as you can get it.

OpenStudy (anonymous):

Thank you. That's the answer i was looking for.

OpenStudy (chriss):

You're welcome